OpenStudy (anonymous):

solve using the zero product property. 3x^3=-7x^2+6x

OpenStudy (anonymous):

Set equal to zero 3x^3+7x^2-6x. Factor out x, x (3x^2 + 7x - 6) x (3x- 2) (x + 3) x = 0 or 3x - 2 = 0 or x + 3 = 0 x = 0, 2/3, -3